(a) Use Euclidean algorithm to find the gcd of 105 and 231.


(b) Use mathematical induction to show that


1 · 1! + 2 · 2! + · · · + n · n! = (n + 1)! − 1


(c) Show that the relation ∼ defined on R as a ∼ b if b−a ∈ Q, is an equivalence relation.


Also, find the equivalence class of 1.


1
Expert's answer
2022-03-09T08:28:37-0500

(a) GCD(105,231)

Set up a division problem where a is larger than b.

a ÷ b = c with remainder R. Do the division. Then replace a with b, replace b with R and repeat the division. Continue the process until R = 0.


231 ÷ 105 = 2 R 21    (231 = 2 × 105 + 21)


105 ÷ 21 = 5 R 0    (105 = 5 × 21 + 0)


When remainder R = 0, the GCF is the divisor, b, in the last equation. GCD = 21 [Answer]


"\sum_{n=-\infty}^{+\infty} f(x) \\"


(b) Use mathematical induction to show that


1 · 1! + 2 · 2! + · · · + n · n! = (n + 1)! − 1

S1=1.1!=1

=(1+1)!-1=2!-1=2-1=1

Sk= 1.1! + 2 · 2! + · · · + k · k! = (k + 1)! − 1

Sk+1= 1.1! + 2 · 2! + · · · + k+1 · (k+1)! = (k+1 + 1)! − 1=(k+2)!-1


Therefore,

Sk+ Ak+1 =Sk+1, where Ak+1 = k+1 · (k+1)!

(k+1)!-1 +k+1 · (k+1)! = (k+1)![1+(k+1)]−1 = (k+1)![k+2]−1=(k+2)!−1 [Answer]


(c) Show that the relation ∼ defined on R as a ∼ b if b−a ∈ Q, is an equivalence relation.


Also, find the equivalence class of 1.



Let us show that a relation '"\sim"' on "\mathbb R" by '"a \sim b" if "a - b"  ∈ Q is an equivalence relation, that is reflexive, symmetric, and transitive.


For each "a\in\mathbb R""a-a=0\in Q" , and thus "a\sim a" each "a\in\mathbb R", and the relation is reflexive.


If "a\sim b", then "a-b\in Q". It follows that "b-a=-(a-b)\in Q", and thus "b\sim a." Consequently, the relation is symmetric.


Let "a\sim b" and "b\sim c". Then "a-b\in Q, \ b-c\in Q". It follows that "a-c=(a-b)+(b-c)\in Q", and thus "a\sim c". Therefore, the relation is transitive.


By definition, the equivalence class of an element "a" is "[a]=\{b\in\mathbb R\ |\ b-a\in Q\}".


Let us give the equivalence classes of 1


"[1]=\{b\in\mathbb R\ |\ b-1\in Q\}\\
=\{b\in\mathbb R\ |\ b=1+n,\ n\in Q\}\\=1+Q" [Answer]